One of the equal sides of an isoceles triangle is 3m less than twice its base. The perimeter is 44. Find the lengths of the sides
Answer by checkley71(8403)

44=(2B-3)+(2B-3)+B
44=5B-6
5B=44+6
5B=50
B=50/5
B=10 ANSWER FOR THE BASE.
SIDE=2*10-3
20-3=17
PROOF
44=17+17+10
44=44